please help me if any one can.....
|1-x|=|2x+3| please help me in solving it with describtion....

Answer by rothauserc(4718)

we are given |1-x| = |2x+3|, there are two equations to consider
1) 1-x = 2x+3
2) -(1-x) = -(2x+3)
3) -(1-x) = 2x+3
1) 1-x = 2x+3
add x to both sides of =
1 = 3x + 3
x = -2/3
2) -(1-x) = -(2x+3)
-1 +x = -2x-3
add 1 to both sides of =
x = -2x -2
add 2x to both sides of =
3x = -2
x = -2/3
3) -(1-x) = 2x+3
-1 +x = 2x+3
add 1 to both sides of =
x = 2x +4
subtract x from both sides of =
-x = 4
x = -4
answer is x = -4, -2/3